OpenStudy (anonymous):

What is the slope of the line through the points (-2, 7) and (3, -3)? m = _?_ (Write only the value of the slope in the answer blank, do not use any variables.)

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (-3-7)/(3-(-2)) m = (-3-7)/(3+2) m = (-10)/5 m = -2 So the slope of the line that passes through the two points is -2

OpenStudy (anonymous):

Thank You

jimthompson5910 (jim_thompson5910):

yw